# Identified-and-Targeted: The First Early Evidence of the Privacy-Invasive Use of Browser Fingerprinting for Online Tracking

### v1

@@ Abstract
To understand the privacy-invasive use of fingerprinting for user tracking, this paper introduces a new framework “FPTrace” (fingerprinting-based tracking assessment and comprehensive evaluation framework) designed to identify alterations in advertisements resulting from adjustments in browser fingerprinting settings. Our approach involves emulating genuine user interactions, capturing advertiser bid data, and closely monitoring HTTP information. Using FPTrace, we conduct a large scale measurement study to identify whether browser fingerprinting is being used for the purpose of user tracking and ad targeting. The results we have obtained provide robust evidence supporting the utilization of browser fingerprinting for the purposes of advertisement tracking and targeting. This is substantiated by significant disparities in bid values and a reduction in HTTP records subsequent to changes in fingerprinting. We additionally demonstrate the potential use of fingerprinting for privacy-evading online tracking purposes even when users opt out of tracking under GDPR/CCPA regulations. In conclusion, our research unveils the widespread employment of browser fingerprinting in online advertising, prompting critical considerations regarding user privacy and data security within the digital advertising landscape.
1 Introduction
@@
Browser fingerprinting is a technique employed to surreptitiously collect data regarding a user’s web browser settings during their online activities. The collected data is then utilized to construct a unique digital identity, commonly referred to as a ‘fingerprint,’ for that specific user browser. Each time a user visits a website, there is potential for the site to employ browser fingerprinting as a means to identify or track the user. Many earlier research studies, such as Englehardt et al. [24] and Lerner et al. [31] and reports such as [1], assumed that the adoption of a fingerprinting script itself is an indication of web tracking and a violation of web privacy. However, this assumption does not hold — just like cookies, browser fingerprinting can be used for defensive security purposes, like bot/fraud detection or authentication. For example, Wu et al. [45] show that adversarial fingerprints are different from those of benign users and therefore many real-world websites are using fingerprints for bot/fraud detection. As another example, Lin et al. [32] have demonstrated the real-world usage of browser fingerprinting in authentication just like what Laperdrix et al. [28] have illustrated in a feasibility study.
@@
Therefore, the research question that we are answering in this paper is: whether browser fingerprints are indeed adopted for online tracking, thus violating web privacy. To the best of our knowledge, none of the prior works have established the link between browser fingerprinting and online tracking. On one hand, many works [24, 31], as mentioned above, consider the existence of fingerprinting scripts as a means of online tracking, which is not true. On the other hand, people have studied the relationship between personalized advertisements and web tracking in general, like cookie-based tracking. For instance, Wills et al. [44] explored ad tracking on the Google and Facebook advertising platforms. Similarly, Zeng et al. [47] employed header bidding to assess targeted ads. These studies did not specifically address the methods employed to link tracking with online advertising; therefore, it remains unclear whether browser fingerprinting contributes to online tracking and thus the privacy violation.
@@
Paper Organization: The remainder of this paper is structured as follows. Section 2 provides an overview of browser fingerprinting and spoofing methodology. Section 3 details the design of FPTrace, while Section 4 outlines our experimental methodology. Section 5 introduces the configuration of experiments including setting details. Section 6 presents our research findings. Section 7 presents the limitation of our research. Section 8 introduces some previous work focusing on the connection between online tracking and targeting. Finally, Section 9 offers our main conclusions and insights.
2 Background
2.1 Browser Fingerprinting
@@
Browser fingerprinting is a technique to identify users using a bunch of data collected from users’ browsers, which may be unique to each browser (we refer to [29] for a representative survey in browser fingerprinting).
@@
Yen et al. [46] and Nikiforakis et al. [39] assess the efficacy of fingerprinting techniques. Vastel et al. [43] examine the discrepancies present in browser fingerprints. Boda et al. [19] and Cao et al. [20] investigate alternative aspects of browser fingerprinting. Previous studies have explored various facets of browser fingerprinting, including techniques involving canvas [36], JavaScript [37], and hardware [38] methods. Information like User Agent which contains the users’ browser version and operating system version can be obtained by the website via HTTP header. Screen resolution, canvas fingerprints, users’ current time zone information can be accessed by the website via JavaScript APIs. Such information can be gathered and used to construct a user profile. Eckersley [23], Fifield et al. [25] and Laperdrix et al. [30] have shown that the browser fingerprints can be used for user identification due to their uniqueness. Acar et al. [18] and Iqbal et al. [27] demonstrated that browser fingerprints can be utilized in fraud detection. In real-world applications, services like Datadome [4], Radware [5], and HUMAN [12] employ browser fingerprinting for the purpose of fraud detection. Previous research such as [41] and articles like [1] solely assessed the presence of fingerprinting-related scripts on websites. However, this alone does not indicate malicious intent for user tracking.
@@
Website “COVER YOUR TRACK” can display each browser fingerprint feature [3]. Website “AM I UNIQUE” can also show the uniqueness of each browser fingerprint among all other users [9]. Website “FingerprintJS” generate a unique ID for each user by using the browser fingerprints, and also display the ID to user [17].
2.2 Browser Fingerprint Spoofing
Spoofing browser fingerprints enables the concurrent operation of multiple browser instances on a single device, each with distinct browser fingerprinting. The Gummy Browser poses a significant threat to the privacy and security of applications that rely on browser fingerprinting [34]. This tool employs three distinct methods of attack, script injection, script modification, and the exploitation of the browser’s inherent settings and debugging tools.
@@ 2.2 Browser Fingerprint Spoofing
Notably, the Gummy Browser exhibits the capability to successfully emulate all JavaScript-based browser fingerprinting techniques, including the sophisticated canvas fingerprinting method. In the context of script modification, the attacker wields the ability to manipulate the behavior of the JavaScript API “toDataURL()” by substituting it with either a predefined or randomized string. Conversely, in the case of script injection, the attacker can strategically introduce a breakpoint within the script precisely at the juncture where the JavaScript API “toDataURL()” is invoked, thereby enabling the replacement of its value with either a predetermined or random string.
2.3 Cookie Restoration
@@
2.4 Header Bidding
Header bidding [6] is a method employed by publishers on websites. Here, publishers designate specific advertising spaces for potential advertisers. The advertiser securing the highest bid gains the chance to display their ads in the corresponding slots. In client-side header bidding, users have the convenience of directly accessing and observing all the bids from their web browsers. Prebid.js [11] is a notable implementation of header bidding. Through the API pbjs.getBidResponses(), users on the client side can inspect the list of advertisers who engaged in the bidding process to secure the opportunity to display ads during the current user’s visit. In the study outlined in [40], the author observes that profiles classified as “Only category” command prices around 40% higher than those assigned to “New user” profiles. The key finding underscores that advertisers’ bidding behavior is shaped by their prior familiarity with the user, resulting in elevated bid values compared to users for whom advertisers lack previous knowledge. Liu et al. [33] additionally demonstrated that advertisers with knowledge of users through data syncing tend to submit higher bid values in header bidding.
2.5 Privacy Regulations
Recently, the European Union introduced the General Data Protection Regulation (GDPR) [15], and California introduced the California Consumer Privacy Act (CCPA) [16], both possessing the potential to regulate and restrain the online advertising and tracking ecosystem. GDPR requires that online services secure user consent (Articles 4 (11)) prior to processing user data (Article 6 (1) (a)). CCPA requires online services to offer users the ability to opt-out of the sale of their personal data (Section 1798 (a) (1)).
Both GDPR and CCPA require that websites must provide privacy notices containing information and controls for opting in/out of the collection and processing of personal information, which should include browser fingerprints. To comply with these requirements, websites are required to integrate consent management platforms (CMPs) [14]. CMPs scan websites, identifying all cookies set by HTTP headers and scripts, no matter first or third-party resources. In the case of GDPR, CMPs must ensure that only strictly necessary cookies are shared, and consent is obtained before sharing non-essential cookies. In the case of CCPA, CMPs should ensure that they provide users with controls to opt-out of the sale of their personal information.
3 FPTrace: Our Framework to Assess Fingerprinting Use in Tracking
In this section, we introduce the design and implementation of our FPTrace framework.
3.1 Framework Design
@@
FPTrace is capable of emulating human user behavior by systematically visiting a curated list of websites categorized under specific keywords, such as “Computer". During the visit, FPTrace not only loads the website but also simulates the behavior of a real human user by browsing the webpage. It is capable of clicking on various links or advertisements, moving the mouse to specific areas on the webpage, and even interacting with functions provided by third-party services, such as Consent Management Platforms (CMPs) or standard cookie banners. After exploring the websites, FPTrace is able to compile a browser profile, encompassing details like browsing history and cookie data. This comprehensive profile is designated as the “Interest Persona”, serving as a representation of user interests and preferences. Furthermore, it has the capability to extract data directly from advertisers, facilitating an in-depth analysis of targeted advertising triggers. FPTrace is capable of capturing and recording data such as bidding behaviors from various advertisers, as well as personal data syncing events. This information is then securely stored within the local database of FPTrace.
@@
Figure 1: High level overview of measurement study methodology. In Step 1, We create browser persona by visiting a list of websites. This step is called “Simulating Interest Personas”. In Step 2, we first use trained persona to visit websites which display ads, then collect bids and HTTP data. This step is called “Collecting Bids and HTTP Data”. In Step 3, we extract cookies from the browser profile and compare them between different experiment settings. This step is called “Detecting Cookie Restoration”. In Step 4, we analyze the manual inspect extracted cookies. This step is called “Detecting Fingerprinting Related Cookie Restoration”.
3.2 Framework Implementation
@@
3.2.1 Web Crawling
We used OpenWPM [24] to construct our crawling FPTrace. OpenWPM is a widely used tool to collect large amount of data related to privacy. OpenWPM relied on the Selenium [13] to handle browser Firefox to do automated crawling.
@@ 3.2.1 Web Crawling
For collecting bids, we used the Prebid.js, a popular implementation of the header bidding protocol [11], mostly because it takes place on the client side, enabling us to intercept the bidding process [2]. The bids objects we get from Prebid.js include all partipated advertisers information, and the ads image HTML contents.
@@
3.2.2 Spoofing Browser Fingerprints
@@
By spoofing JavaScript APIs, we used real browser fingerprints data from an dataset on Github [42]. The data is collected from real devices. In spoofing JavaScript APIs, we used Script Injection code provided by Gummy Browser [35] to overwrite all possible JavaScript APIs that can be used in browser fingerprints, such as Navigator(API: Navigator), Screen(API: Screen), Canvas(API: canvas.getContext) and Date(API: Date) APIs. To ensure the spoofing of all APIs prior to the loading of any webpage, we encapsulated all spoofing scripts within a browser extension.
Only spoofing JavaScript APIs may not work if the websites match the value from JavaScripts to the one from HTTP header. So we have also spoof the HTTP header value. ModHeader [8] is an extension to allow users to modify HTTP headers. We embedded Selenium ModHeader [7] version into OpenWPM to spoof value in HTTP header.
3.2.3 Capturing Bidding Object
Header bidding can signal the enthusiasm of potential bidders. A bidder interested in showcasing their advertisements in the present slot will typically place higher bids. Cook et al. [22] have proved that the header bidding can be used to identify if the user is tracked. A higher bid can represent the user is tracked, and the user information is used by publishers, bidders, or bidding auctions. Specifically, we use Prebid.js to monitor bidders’ behaviors. Prebid.js is an implementation of header bidding protocol, and the user can get bidders’ behaviors, including the winner of each ad slot, the ads image HTML contents, and also all of the participated bidders bids value no matter the bidders win or lose.
@@ 3.2.4 Exporting Cookies
4 Measurement Study Methodology
In this section, our measurement study methodology that utilizes the FPTrace framework is introduced. Initially, in Section 4.1, we outline the simulation of interest personas, which emulate real person’s interests. Following this, in Section 4.2, the collection of bids and HTTP data using various settings is presented, employing simulated interest personas to visit websites that use Prebid.js. Subsequently, the detection of cookie restoration post removal is discussed in Section 4.3, followed by an elaboration on the identification of browser fingerprinting-associated cookie restoration in Section 4.4. Figure 1 provides an overview of the measurement study methodology. Finally, in Section 4.5, the configuration for discerning whether browser fingerprinting persists for data collection post-user opt-out under GDPR and CCPA protection is discussed.
4.1 Simulating Interest Personas
@@
Figure 2: High level overview of advertisement experiment. In step 1, FPTrace will visit websites sequentially to keep updating the browser profile. In Step 2, FPTrace will control the browser with updated profile to visit each website. In Step 3, FPTrace will record the bids data and HTTP data. In Step 4, FPTrace will not update the current browser profile, and load the profile updated after Step 1. After visiting all websites in Step 2, FPTrace will compare the bids data and HTTP data between A, A* and B.
4.2 Collecting Bids and HTTP Data

### v2

# The First Early Evidence of the Use of Browser Fingerprinting for Online Tracking

@@ Abstract.
This paper introduces “FPTrace” (fingerprinting-based tracking assessment and comprehensive evaluation framework), a framework to assess fingerprinting-based user tracking by analyzing ad changes from browser fingerprinting adjustments. Using FPTrace, we emulate user interactions, capture ad bid data, and monitor HTTP traffic. Our large-scale study reveals strong evidence of browser fingerprinting for ad tracking and targeting, shown by bid value disparities and reduced HTTP records after fingerprinting changes. We also show fingerprinting can bypass GDPR/CCPA opt-outs, enabling privacy-invasive tracking.
In conclusion, our research unveils the widespread employment of browser fingerprinting in online advertising, prompting critical considerations regarding user privacy and data security within the digital advertising landscape.
copyright: none
1. Introduction
@@
Browser fingerprinting is a technique employed to surreptitiously collect data regarding a user’s web browser settings during their online activities. The collected data is then utilized to construct a unique digital identity, commonly referred to as a ‘fingerprint,’ for that specific user browser. Each time a user visits a website, there is potential for the site to employ browser fingerprinting as a means to identify or track the user. Many earlier research studies, such as Englehardt et al. (Englehardt and Narayanan, 2016) and Lerner et al. (Lerner et al., 2016) and reports such as (inc, 2023), assumed that the adoption of a fingerprinting script itself is an indication of web tracking and a violation of web privacy. However, this assumption does not hold — just like cookies, browser fingerprinting can be used for defensive security purposes, like bot/fraud detection or authentication. For example, Wu et al. (Wu et al., 2023) show that adversarial fingerprints are different from those of benign users and therefore many real-world websites are using fingerprints for bot/fraud detection. As another example, Lin et al. (Lin et al., 2022) have demonstrated the real-world usage of browser fingerprinting in authentication just like what Laperdrix et al. (Laperdrix et al., 2019) have illustrated in a feasibility study.
@@
Therefore, the research question that we are answering in this paper is: whether browser fingerprints are indeed adopted for online tracking, thus violating web privacy. To the best of our knowledge, none of the prior works have established the link between browser fingerprinting and online tracking. On one hand, many works (Englehardt and Narayanan, 2016; Lerner et al., 2016), as mentioned above, consider the existence of fingerprinting scripts as a means of online tracking, which is not true. On the other hand, people have studied the relationship between personalized advertisements and web tracking in general, like cookie-based tracking. For instance, Wills et al. (Wills and Tatar, 2012) explored ad tracking on the Google and Facebook advertising platforms. Similarly, Zeng et al. (Zeng et al., 2022) employed header bidding to assess targeted ads. These studies did not specifically address the methods employed to link tracking with online advertising; therefore, it remains unclear whether browser fingerprinting contributes to online tracking and privacy violation.
@@
2. Background
2.1. Browser Fingerprinting
@@
Browser fingerprinting is a technique to identify users using data collected from browsers, which may be unique to each browser (we refer to (Laperdrix et al., 2020) for a representative survey in browser fingerprinting).
@@
Yen et al. (Yen et al., 2012) and Nikiforakis et al. (Nikiforakis et al., 2013) assess the efficacy of fingerprinting techniques. Vastel et al. (Vastel et al., 2018) examine the discrepancies present in browser fingerprints. Boda et al. (Boda et al., 2012) and Cao et al. (Cao et al., 2017) investigate alternative aspects of browser fingerprinting. Previous studies have explored various facets of browser fingerprinting, including techniques involving canvas (Mowery and Shacham, 2012), JavaScript (Mulazzani et al., 2013), and hardware (Nakibly et al., 2015) methods. Information like User Agent which contains the users’ browser version and operating system version can be obtained by the website via HTTP header. Screen resolution, canvas fingerprints, users’ current time zone information can be accessed by the website via JavaScript APIs. Such information can be gathered and used to construct a user profile. Eckersley (Eckersley, 2010), Fifield et al. (Fifield and Egelman, 2015) and Laperdrix et al. (Laperdrix et al., 2016) have shown that the browser fingerprints can be used for user identification due to their uniqueness. Acar et al. (Acar et al., 2013) and Iqbal et al. (Iqbal et al., 2021) demonstrated that browser fingerprints can be utilized in fraud detection. In real-world applications, services like Datadome (dat, 2023), Radware (rad, 2023), and HUMAN (hum, 2023) employ browser fingerprinting for the purpose of fraud detection. Previous research such as (Sjösten et al., 2021) and articles like (inc, 2023) solely assessed the presence of fingerprinting-related scripts on websites. However, this alone does not indicate malicious intent for user tracking.
@@
Website “COVER YOUR TRACK” can display each browser fingerprint feature (cov, 2023). Website “AM I UNIQUE” can also show the uniqueness of each browser fingerprint among all other users (ami, 2023). Website “FingerprintJS” generate a unique ID for each user by using the browser fingerprints, and also display the ID to user (fin, 2020).
2.2. Browser Fingerprint Spoofing
@@
Spoofing browser fingerprints enables the concurrent operation of multiple browser instances on a single device, each with distinct browser fingerprinting. The Gummy Browser poses a significant threat to the privacy and security of applications that rely on browser fingerprinting (Liu et al., 2022). This tool employs three distinct methods of attack, script injection, script modification, and the exploitation of the browser’s inherent settings and debugging tools. Notably, the Gummy Browser exhibits the capability to successfully emulate all JavaScript-based browser fingerprinting techniques, including the sophisticated canvas fingerprinting method. In script modification, the attacker wields the ability to manipulate the behavior of the JavaScript API “toDataURL()” by substituting it with either a predefined or randomized string. Conversely, in the case of script injection, the attacker can strategically introduce a breakpoint within the script precisely at the juncture where the JavaScript API “toDataURL()” is invoked, thereby enabling the replacement of its value with either a predetermined or random string.
2.3. Cookie Restoration
@@
2.4. Header Bidding
Header bidding (HB_, 2023) is a method employed by publishers on websites. Here, publishers designate specific advertising spaces for potential advertisers. The advertiser securing the highest bid gains the chance to display their ads in the corresponding slots. In client-side header bidding, users have the convenience of directly accessing and observing all the bids from their web browsers. Prebid.js (pre, 2023) is a notable implementation of header bidding. Through the API pbjs.getBidResponses(), users on the client side can inspect the list of advertisers who engaged in bidding process to secure the opportunity to display ads during the current user’s visit. In the study outlined in (Olejnik et al., 2014), the author observes that profiles classified as “Only category” command prices around 40% higher than those assigned to “New user” profiles. The key finding underscores that advertisers’ bidding behavior is shaped by their prior familiarity with the user, resulting in elevated bid values compared to users for whom advertisers lack previous knowledge. Liu et al. (Liu et al., 2024) additionally demonstrated that advertisers with knowledge of users through data syncing tend to submit higher bid values in header bidding.
2.5. Privacy Regulations
Recently, the European Union introduced the General Data Protection Regulation (GDPR) (GDP, 2016), and California introduced the California Consumer Privacy Act (CCPA) (CCP, 2018), both possessing the potential to regulate and restrain the online advertising and tracking ecosystem. GDPR requires that online services secure user consent (Articles 4 (11)) prior to processing user data (Article 6 (1) (a)). CCPA requires online services to offer users the ability to opt-out of the sale of their personal data (Section 1798 (a) (1)). Both require websites to provide privacy notices containing information and controls for opting in/out of the collection and processing of personal information, which should include browser fingerprints. To comply with these requirements, websites are required to integrate consent management platforms (CMPs) (Web, 2023). CMPs scan websites, identifying all cookies set by HTTP headers and scripts from first and third-party resources. In the case of GDPR, CMPs must ensure that only necessary cookies are shared, and consent is obtained before sharing non-essential cookies. In the case of CCPA, CMPs should ensure that they provide users with controls to opt-out of the sale of personal information.
3. FPTrace: Our Framework to Assess Fingerprinting Use in Tracking
3.1. Framework Design
@@
FPTrace is capable of emulating human user behavior by systematically visiting a curated list of websites categorized under specific keywords, such as “Computer”. During the visit, FPTrace not only loads the website but also simulates the behavior of a real human user by browsing the webpage. It is capable of clicking on various links or advertisements, moving the mouse to specific areas on the webpage, and even interacting with functions provided by third-party services, such as Consent Management Platforms (CMPs) or standard cookie banners. After exploring the websites, FPTrace is able to compile a browser profile, encompassing details like browsing history and cookie data. This comprehensive profile is designated as the “Interest Persona”, serving as a representation of user interests and preferences. Furthermore, it has the capability to extract data directly from advertisers, facilitating an in-depth analysis of targeted advertising triggers. FPTrace is capable of capturing and recording data such as bidding behaviors from various advertisers, as well as personal data syncing events. This information is then securely stored within the local database of FPTrace.
@@
3.2. Framework Implementation
@@
3.2.1. Web Crawling
We used OpenWPM (Englehardt and Narayanan, 2016) to construct our crawling FPTrace. OpenWPM is a widely used tool to collect large amount of data related to privacy. OpenWPM relied on the Selenium (Sel, 2023b) to handle browser Firefox to do automated crawling.
@@ 3.2.1. Web Crawling
For collecting bids, we used the Prebid.js, a popular implementation of the header bidding protocol (pre, 2023), mostly because it takes place on the client side, enabling us to intercept the bidding process (cli, 2023). The bids objects we get from Prebid.js include all partipated advertisers information, and the ads image HTML contents.
@@
3.2.2. Spoofing Browser Fingerprints
@@
By spoofing JavaScript APIs, we used real browser fingerprints data from an dataset on Github (Spirals-Team, 2020). The data is collected from real devices. In spoofing JavaScript APIs, we used Script Injection code provided by Gummy Browser (Liu et al., 2022) to overwrite all possible JavaScript APIs that can be used in browser fingerprints, such as Navigator(API: Navigator), Screen(API: Screen), Canvas(API: canvas.getContext) and Date(API: Date) APIs. To ensure the spoofing of all APIs prior to the loading of any webpage, we encapsulated all spoofing scripts within a browser extension.
Only spoofing JavaScript APIs may not work if websites match the value from JavaScripts to the one from HTTP header. So we have also spoofed the HTTP header. ModHeader (Mod, 2023) is an extension to allow users to modify HTTP headers. We embedded Selenium ModHeader (Sel, 2023a) version into OpenWPM to spoof HTTP header.
3.2.3. Capturing Bidding Object
Header bidding can signal the enthusiasm of potential bidders. A bidder interested in showcasing their advertisements in the present slot will typically place higher bids. Cook et al. (Cook et al., 2020) have proved that the header bidding can be used to identify if the user is tracked. A higher bid can represent the user is tracked, and the user information is used by publishers, bidders, or bidding auctions. Specifically, we use Prebid.js to monitor bidders’ behaviors. Prebid.js is an implementation of header bidding protocol, and the user can get bidders’ behaviors, including the winner of each ad slot, the ads image HTML contents, and also all of the participated bidders bids value.
@@ 3.2.4. Exporting Cookies
Figure 1. High level overview of measurement study methodology. In Step 1, We create browser persona by visiting a list of websites. This step is called “Simulating Interest Personas”. In Step 2, we first use trained persona to visit websites which display ads, then collect bids and HTTP data. This step is called “Collecting Bids and HTTP Data”. In Step 3, we extract cookies from the browser profile and compare them between different experiment settings. This step is called “Detecting Cookie Restoration”. In Step 4, we analyze the manual inspect extracted cookies. This step is called “Detecting Fingerprinting Related Cookie Restoration”.
4. Measurement Study Methodology
In this section, our measurement study methodology that utilizes the FPTrace framework is introduced. Initially, in Section 4.1, we outline the simulation of interest personas, which emulate real person’s interests. Following this, in Section 4.2, the collection of bids and HTTP data using various settings is presented, employing simulated interest personas to visit websites that use Prebid.js. Subsequently, the detection of cookie restoration post removal is discussed in Section 4.3, followed by an elaboration on the identification of browser fingerprinting-associated cookie restoration in Section 4.4. Figure 1 provides an overview of the measurement study methodology. Finally, in Appendix A.3, the configuration for discerning whether browser fingerprinting persists for data collection post-user opt-out under GDPR and CCPA protection is discussed.
4.1. Simulating Interest Personas
@@
Figure 2. High level overview of advertisement experiment. In step 1, FPTrace will visit websites sequentially to keep updating the browser profile. In Step 2, FPTrace will control the browser with updated profile to visit each website. In Step 3, FPTrace will record the bids data and HTTP data. In Step 4, FPTrace will not update the current browser profile, and load the profile updated after Step 1. After visiting all websites in Step 2, FPTrace will compare the bids data and HTTP data between A, A* and B.
4.2. Collecting Bids and HTTP Data